OpenStudy (anonymous):

A heater dissipates a power of 400 W and the potential difference across it is 110 V. If the potential difference is changed to 220 V keeping its resistance constant, what will be its power dissipation? 400 W 800 W 1600 W cannot be determined

OpenStudy (anonymous):

the relation between power , voltage and resistance is P=V² / R using this relation in both the cases P = V² / R - (1) P' =V'² / R' - (2) R=R' - given dividing (1) and (2) we get P / P' = V² / V'² inserting the values we find 400 / P' = 110² / 220² therefore P' = 800 W

OpenStudy (anonymous):

Use the formula: P = V^2 / R. P1 = 400 W and V1 = 110 V Since P = V^2 / R, 400 = (110^2) / R, and R = 30.25 ohms If V2 = 220 V then P = (220^2) / 30.25 and P = 1600 W Another method: R1 = (V1^2) / P1 and R2 = (V2^2) / P2 where V1 = 110 V, P1 = 400 W, V2 = 220 V, and P2 is what you're trying to find. Since R remains constant, R1 = R2 and (V1^2) / P1 = (V2^2) / P2. (110^2) / 400 = (220^2) / P2 so P2 = 1600 W. ANSWER: P = 1600 W.

OpenStudy (anonymous):

sorry the answer is P' = 1600 calculation error